OpenStudy (anonymous):

Solve the following system by hand. List final answers as ordered pairs. x - y + z = -6 4x - y + 3z = -11 -5x + 2y - z = 12

OpenStudy (anonymous):

eliminate one variable

OpenStudy (anonymous):

let's eliminate z x-y+z=-6 z=-6-x+y

OpenStudy (anonymous):

x-y+z=-6 y=x+z+6

OpenStudy (anonymous):

4x-y+3z=-11 y=4x+3z+11

OpenStudy (anonymous):

-5x+2y-z=12 y=5x+z+12/2